OpenStudy (anonymous):

how do you find the x-intercept of a line if you only have the slope(2/3) and y-intercept (0,5)

OpenStudy (raden):

first u have to make an equation of the line is y=mx+c, for this case given m=2/3 and c=5. so, it is y=(2/3)x+5 now, to find x-intercept, take y=0, then solve for x